OpenStudy (anonymous):

Could someone tell me the formulas for converting radians to degrees and degrees to radians? Thanks!

hartnn (hartnn):

sure.!

OpenStudy (anonymous):

use unitary method \[\Pi=180 degree\]

hartnn (hartnn):

remember \(\pi ^r = 180 ^o\) so, \(x^o = \pi/180 \) radians and \(x^r = 180/ \pi\)degrees.

OpenStudy (anonymous):

Thank you guys! (:

OpenStudy (anonymous):

to convert degrees into radion multiply the degree with \[\Large \frac{\pi}{180}\] to convert radion into degrees multiply he radians with \[\Large \frac{180}{\pi}\]

hartnn (hartnn):

sorry, mistyped it \(x^o = x \times \pi/180\) and \(x^r= x \times 180/\pi \)

hartnn (hartnn):

want examples ?

OpenStudy (anonymous):

Sure o: This one problem I'm doing now is converting 11pi/12 to degrees.

OpenStudy (anonymous):

And thanks everyone for contributing (:

hartnn (hartnn):

so what you'll multiply that with ? 11pi/2 * .... ?

hartnn (hartnn):

180/pi or pi/180 ??

OpenStudy (anonymous):

180/pi.. I think you switched them back up there o:

hartnn (hartnn):

no, i didn't switch and 180/pi is correct, so what you got as degree angle ?

OpenStudy (anonymous):

you wrote to get radians, you use 180/pi. o: and 165.

hartnn (hartnn):

ohh!!

OpenStudy (anonymous):

Lol aw no biggie, I appreciate the help (:

hartnn (hartnn):

sorry, my bad. \(x^o \times \pi/180=x^r \\x^r \times 180/\pi=x^o\) and 11pi/2 * 180/pi = 11 *180 /2 = 11*90 = 990 how u got 165 ?

OpenStudy (anonymous):

11pi/12 o: Lool.

hartnn (hartnn):

i need some sleep xD

hartnn (hartnn):

165 is corret.

hartnn (hartnn):

correct* damn

OpenStudy (anonymous):

Lool, get to bed! (: Thanks again!

hartnn (hartnn):

welcome ^_^